Q10.

For a given matrix A = "\begin{bmatrix}
 1 & 0 & 2 & -3 \\
 2 & 0 & 4 & - 6 \\
 -3 & 0 & - 6& 9
\end{bmatrix}". Which of the following is true

(1) rank (A) = 3, nullity (A) = 1

(2) rank (A) = 2, nullity (A) = 2

(3) rank (A) = 1, nullity (A) = 3

(4) None of the given answers is true.


Q11.

For a, b "\in" R, the transformation T : R^2 "\to" R^3 defined byT (x, y) = (2x - y, 3x + y + 3A, 5x - 2y + bxy) is linear if

(1) a = b = 1

(2) a = 0; b = 1

(3) a = 1; b = 0

(4) None of the given answers is true.


Q12.

Suppose T : R^3 "\to" R^2

is a linear defined by

T (x, y) = (4x + 3y + Z, x - 2y). Then which of the following is basis of range T is

(1) (4, 1, 3), (1, 2, 0)

(2) (1, 0), (0, 1), (1, 1)

(3) (4, 1), (3, 2), (1, 0)

(4) None of the given answers is true.


1
Expert's answer
2022-01-19T18:02:39-0500

Q 10

Reduced row Echolon form of A


"rref\>A=\begin{pmatrix}
 1&0&2&-3 \\
 0&0&0&0 \\
0&0&0&0
\end{pmatrix}"


Rank (A)= 1, Nullity (A) =3


Q 11


If T is linear it should satisfy

T(cx,cy)=cT(x,y)

For any scalar c, and any element x,y in the domain of T.


T(cx,cy)= (2cx-cy, 3cx+cy+3a, 5cx-2cy+bcxcy)


"\ne" c(2x-y, 3x+y+3a, 5x-2y+bxy)


None of the given answer is true.


Q 12


"\begin{pmatrix}
 4x+3y+z \\
 x-2y
\end{pmatrix}=x\begin{pmatrix}
 4 \\
 1 
\end{pmatrix}" "+y\begin{pmatrix}
 3 \\
 -2
\end{pmatrix}+z\begin{pmatrix}
 1 \\
 0
\end{pmatrix}"


Transforming matrix


"A=\begin{pmatrix}
 4&3&1 \\
 1&-2 & 0\\

\end{pmatrix}"



rref A "=\begin{pmatrix}
 1&0&\frac{2}{11} \\
 0&1&\frac{1}{11} 
\end{pmatrix}"


Basis for range T is (1,0),(0,1)


None of the given answers is true.